# {strip}[Uncaptioned image]

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
To address this issue, and as part of our broader project addressing NCIM, we developed Hands-Off, an interaction technique for messaging applications that deters non-consensual screenshots (see Fig. 1). After a formative study, we built Hands-Off to require recipients to perform a hand gesture in the air, above the device, to unlock media. Borrowing from the social meaning of gestures (e.g., handshakes) (Müller, 2004; Zhou and Zhang, 2008), Hands-Off uses gestures as design friction to convey a user’s consent preferences, and acts as a deterrent against screenshots. We built a set of four hand-gesture-based deterrence features. If a sender wishes to share a photo, but only wants the receiver to have controlled, momentary access to it, they can require that the receiver perform a gesture. The media will only reveal itself while the hand-gesture is being performed. We also describe a threat model and design space for how gestures can be used with different content types, relationships, and locations.
@@ 2.2. Visual privacy
It has never been easier to take and share photos online. While photo-sharing offers social value to society (e.g. sending a baby photo to grandparents), photo-sharing via wearable devices, ubiquitous surveillance, and facial recognition technologies creates privacy risks. Privacy-inspired systems addresses this concern by focusing primarily in three areas: physical photography, online photo access, and photo recapture. Prior work includes systems to prevent unwanted photography in public spaces, such as Privacy.tag’s QR codes that can be placed on clothing to enact privacy protections (Bo et al., 2014) or using smart LED lighting design in physical spaces to prevent photography (Zhu et al., 2017). The second area of research focuses on limiting users’ access to photos. This includes asking individuals depicted in a photo for preferences regarding obfuscation (Olteanu et al., 2018; Shu et al., 2017, 2018) and drawing cartoons over video streams to protect bystanders’ privacy (Hassan et al., 2017), or only showing the faces to friends on social media (Ilia et al., 2015). A final thread of research includes systems to prevent recapture of a mobile screen, such as preventing second screen capture by changing the lighting and display on a screen (Zhang et al., 2015) and digital watermarking systems that collect metadata of the device making screenshots (Gu et al., 2023; Bai et al., 2023; Jiang et al., [n. d.]).
@@ 2.3. Design friction
Design friction has been implemented in social media platforms to avoid unsafe or undesired actions, particularly ones that have downstream harms on other users. Wang et. al deployed a countdown timer on Facebook to ensure the user waits 20 seconds before making a post in order to minimize regrettable posts (Wang et al., 2014). Gould et. al al notes several design friction choices on Twitter’s platform to warn users against potential misinformation in some posts, as well as asking users to confirm they have read an article themselves before re-tweeting (Gould et al., 2021).
Friction can also be a tool towards meaningful reflection and intention. It is common for functionally opposing design approaches to remove key aspects of their function in order to highlight their values. This includes works in critical design, counter-functional design, slow design, reflective design, and lucid design  (Pierce and Paulos, 2014; Sengers et al., 2005; Gaver and Martin, 2000). Physical friction can also be designed to induce user reflection. For example, Crank That is a physical hand-crank intervention to increase friction and awareness of passive consumption in Twitter content feeds (Song et al., 2021). Users are required to continuously turn a hand crank in order to power a screen to use social media, allowing users to “question how they value various interactions on social media by transforming their usage into an embodied one in which they have to physically work for their experience” (Song et al., 2021).
@@
An effective proactive mechanism should prevent or deter unwanted behavior. Currently, only few messaging platforms allow users to prevent screenshots altogether. [^2] https://www.safedigitalintimacy.org/state-of-the-industry StopNCII, a hashing-based system developed by multiple stakeholders including Meta, allows users to upload photos to a server before these photos are ever posted by someone else, and then block subsequent attempts by someone else to upload these photos. [^3] https://stopncii.org/ However, even then there are many drawbacks. Transformed images may evade the hash algorithm. When these images continue to be reposted, platforms may overly rely on StopNCII and invest less effort in removals. And of course, users must trust the third-party–in this case, Meta. Our work aims to make the burden more equitable, placing “the work” on a potential perpetrator rather than the potential victim.
@@ 3. Formative Study
We conducted a formative interview study to develop design requirements for a technological intervention. The formative study aimed to augment prior qualitative and survey research on NCIM (Citron and Franks, 2014; Citron, 2018, 2022; Franks, 2014; Geeng et al., 2020; Powell et al., 2019; McGlynn et al., 2021; Freed et al., 2017; Qin et al., 2024). The first portion of the interview focused on participants’ existing screenshot experiences and practices. The second portion asked participants for their thoughts about using a screenshot deterrence system. In our conception of Hands-Off, we believed that a gestural interface–one that prevented viewers from using the screenshot features by occupying their hands–could be an effective deterrent. We used the formative study as an opportunity to test this idea. Thus, in the second part of the interview, we had participants explore how they would use a screenshot deterrence feature. We described the design of a gestures-based approach and solicited reflection and feedback feedback. Our university’s IRB approved the study.
@@ 3.1. Participants
Eight participants were recruited from two undergrad courses at a large university. We specifically recruited participants who were able and willing to share their own or other peoples’ experiences about screenshots. We gave participants the option of either 1:1 solo interviews or paired interviews in which the participant brought a friend, who often acted as a confidante, to encourage more conversation. Each interview session took between 45 to 60 minutes. Each participant was compensated \$15 for a pair interview or \$20 for a solo interview.
@@ 3.2.2. Theme 2: Perceived security of gesture-based screenshot prevention
We presented participants with a possible gesture-based approach, using hand gestures to both indicate preferences about and deter from unwanted screenshots. Participants expressed a strong preference for gesture-based security features within mobile messaging applications. Their feedback shed light on the underlying motivations and practical benefits. Gesture-based security provides a tangible barrier, making it significantly harder for the recipient to share content.
@@
Participants noted that gesture-based deterrence could reduce potential awkwardness associated with directly requesting confidentiality through written or spoken messages. This suggests that a gestural approach may enhance privacy while streamlining communication.
@@ 3.2.3. Theme 3: Gesture should depend on context
All participants believed that there should be a variety of gestures to choose from to account for various usage contexts. Specifically, context depends on three axes: the type of content sent, relationship with recipient, and the recipient’s location of viewing. Table 1 maps the connection between the axis and its design considerations.
@@
Table 1. Design considerations from formative study. The type of gesture matters for of content sent, relationship between sender and recipient, and location for viewing the content.
@@ 4.1.1. Threat model
Figure 3. Threat model of possible methods of screen duplication after content is sent. Duplicated content may be posted online repeatedly, continuing this cycle. Hands-Off contributes to deterrence against screenshotting content. Also listed is prior work and where it fits in this threat model.
The threat model for Hands-Off is unwanted photo duplication in online sharing (Bedi et al., 2013) (see Fig. 3). The system should deter from general unauthorized duplication, which would take media out of its original context. We borrow from prior visual privacy systems and “assume that the adversaries are the users, the online services and third parties (e.g., external observers)” (Olteanu et al., 2018). The paradigm begins with a sender who sends a piece of original content to a recipient. From there, the recipient may download the file, screen record, screenshot, capture it using a second device themselves, or utilize an accomplice to capture on a different device.
Hands-Off goals include preventing screenshots, and deterring second device capture by the same individual. Some portions of the threat model are accounted for in prior work. In-app downloads and screen recording may be disabled by the system developer, and second device capture may be deterred with a system such as Kaleido that uses lighting patterns to disrupt outside cameras (Zhang et al., 2015). Screenshots on the device may also be prevented within the operating system of mobile devices, but this is rarely implemented in popular mobile messaging applications. As we have observed, there is no approach that can absolutely guarantee that an image not be captured and reshared. However, it is worth considering which approaches raise the cost of different types of “attack.”
@@ 4.1.2. Design goals
Hands-Off functions by occupying both hands to prevent individuals from engaging the screenshot features on the device. As the hands are occupied, this may further deter second device capture. [^4] Clearly, this is just a deterrent. A particularly motivated attacker could position a second device, record a video, clip frames, and recover a fairly good representation. However, our analysis of themes in the formative study reveal additional specific design constraints:
@@
In addition to the explicit guidelines above, we have opted to focus on abstract gestures with relatively little cultural significance. Some gestures are culturally dependent in meaning (e.g., palms touching in a prayer gesture) whereas others have existing iconic or metaphoric roles (e.g., hands making a heart symbol). By relying on more abstract gestures, we hope that their meaning may be more flexible. This allows some degree of communicable understanding through gestures, while still allowing flexibility for how users choose interpret them. That said, depending on context, gestures with existing meaning may be desirable for certain applications. These are possible with Hands-Off, but are left to future work.
@@ 4.2. Design space
Hands-Off gestures are shaped with two key factors in mind: preventing misuse and fitting within societal norms. Like a literary symbol, social gestures are imbued with meaning. They communicate through physicalization of language or ceremony (Kang and Tversky, 2016; Schiffrin, 1974). A wave goodbye between friends, a handshake after a negotiation, or a raised hand in a court of law are all means of communicating between parties but also to observers. They have semantic meaning (Xu et al., 2009), which may be learned, innate (Iverson and Goldin-Meadow, 1997), or inferred (e.g., hands with palms outward to represent a ‘wall’) (Müller et al., 2013). Even a gesture made in private can embody a sentiment or belief or serve a regulatory function (Rodríguez and Palacios, 2007). Using a gesture, even an abstract one, can physicalize an affirmation, whether observed by someone else, or not: “I understand that the image was shared with me and is not intended for sharing with others.” The friction imposed by the physical nature of the gesture further acts to reinforce the meaning and, thus, a broader value system of consent, autonomy, and agency. It is not a button to be clicked away.
As a form of body language, the choice of gestures in interaction must take into account likely social interpretation. This is key as Hands-Off exists within two overlapping, but distinct, design spaces: bodily friction and social friction. Bodily friction is the invisible but physical restraint. The constrained space is the gap between the user when performing the necessary gesture, and the physical location of the two phone buttons that need to be pressed simultaneously to screenshot. In Hands-Off, the gap is maintained by the system’s detection of a hand gesture and the distance between reaching over and pressing the two buttons. Along the social dimension, someone using their phone in a public space will consider their social appearance of doing so. Prior work has shown that people generally want to avoid making social blunders in the presence of others, and subtle movements and movements that look and feel like everyday actions are preferred over uncommon movements (Rico and Brewster, 2010).
@@ 4.3. Use cases
Silly moment just for us. Sam is having a good time and wants to share a photo of him out drinking with his long distance best friend Ray. However, Sam doesn’t want Ray to send this photo to anyone else. He chooses Hands-Off and selects the binoculars option to match the silliness of the photo with the gesture, and also to imply Ray should not reshare this photo. When Ray receives the message, he is prompted to perform the hand gesture and then sees the photo Sam has shared.
@@ 4.4. Implementation details
We built Hands-Off using SwiftUI for iOS devices. We choose iPhone for it’s consistent button placement and hardware design. We rely on pose detection using a custom CoreML hand pose detection model trained with a custom corpus we built. We defined a multi-class neural-network classification model with 20 convolution layers for interlace, wave, frame, binoculars, and a background class for when none of the above gestures are detected. Training data consisted of 120 to 200 photos of hands performing each gesture per class. We collected the vast majority of the data ourselves by taking photos of hands of participants. We account for high and low lighting, light and dark skin tones, and different photography angles. The remaining photos are copyright-free images of hands from the internet. The model was trained to recognize the hands through the positions of the finger tips and the locations of the joints. See Table 2 for model performance. Facial data was not used to train the model.
@@ 4.5. Hands-Off limitations
One limitation of using the CoreML model is it consistently recognized interlaced hands as one hand instead of two, which impacted precision for interlace (see Table 2). For each other hand gesture, Hands-Off required two hands detected as well as gesture detected to display the photo. However, we had to silently change this two-hands criteria for the interlace option. This limitation tells us the CoreML vision model works well for detected separated hands or poses, but may be limited for gestures when the hands stack or otherwise overlay each other.
We note several limitations in Hands-Off. First, Hands-Off is based on assumptions and modes of interactions specific to the iPhone. Screenshot deterrence is built on top of one static assumption of the double button press on the left and right sides of the phone body. The methods of screenshots can vary in other designs of mobile phones. As future work progresses, limitations may also arise in securing hand gesture-based consent in AR, VR, and XR spaces due to the possibility of certain gestures already being used as input by the specific system. Second, this study tackles just one aspect of many potential privacy concerns in the threat model. Screen recording can still be used to capture the screen if that feature is not disabled in the application. Future research in gesture-based deterrence should consider allowing users to create their own custom hand gestures. This personalization would add a uniquely intimate layer to the interaction, similar to how secret handshakes signify closeness between two people or some in-group status. Finally, Hands-Off is limited in accommodating all abilities (e.g., people with mobility impairments; people who are Blind or low vision). Feminism should acknowledge intersectionality, and that includes building for various abilities. Expanding the range of gesture-based methods (e.g. offering a voice-based alternative), to accommodate users with different abilities, needs, and various device models is crucial.

### v3

@@ Abstract.
Authors emails: rrll,flameiro,shefalip,cisaula,eadar,eegg,yardi@umich.edu
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; October 13–15, 2024; Pittsburgh, PAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
To address this issue, and as part of our broader project addressing NCIM, we developed Hands-Off, an interaction technique for messaging applications that deters non-consensual screenshots (see Fig. 1). After a formative study, we built Hands-Off to require recipients to perform a hand gesture in the air, above the device, to unlock media. Borrowing from the social meaning of gestures (e.g., handshakes) (Müller et al., 2013; Zhou and Zhang, 2008), Hands-Off uses gestures as design friction to convey a user’s consent preferences, and acts as a deterrent against screenshots. We built a set of four hand-gesture-based deterrence features. If a sender wishes to share a photo, but only wants the receiver to have controlled, momentary access to it, they can require that the receiver perform a gesture. The media will only reveal itself while the hand-gesture is being performed. We also describe a threat model and design space for how gestures can be used with different content types, relationships, and locations.
@@ 2.2. Visual privacy
It has never been easier to take and share photos online. While photo-sharing offers social value to society (e.g. sending a baby photo to grandparents), photo-sharing via wearable devices, ubiquitous surveillance, and facial recognition technologies creates privacy risks. Privacy-inspired systems address this concern by focusing primarily in three areas: physical photography, online photo access, and photo recapture. Prior work includes systems to prevent unwanted photography in public spaces, such as Privacy.tag’s QR codes that can be placed on clothing to enact privacy protections (Bo et al., 2014) or using smart LED lighting design in physical spaces to prevent photography (Zhu et al., 2017). The second area of research focuses on limiting users’ access to photos. This includes asking individuals depicted in a photo for preferences regarding obfuscation (Olteanu et al., 2018; Shu et al., 2017, 2018) and drawing cartoons over video streams to protect bystanders’ privacy (Hassan et al., 2017), or only showing the faces to friends on social media (Ilia et al., 2015a). A final thread of research includes systems to prevent recapture of a mobile screen, such as preventing second screen capture by changing the lighting and display on a screen (Zhang et al., 2015) and digital watermarking systems that collect metadata of the device making screenshots (Gu et al., 2023; Bai et al., 2023; Jiang et al., [n. d.]), and photos that fade with time (Pias et al., 2022).
@@ 2.3. Design friction
Design friction has been implemented in social media platforms to avoid unsafe or undesired actions, particularly ones that have downstream harms on other users. Wang et. al deployed a countdown timer on Facebook to ensure the user waits 20 seconds before making a post in order to minimize regrettable posts (Wang et al., 2014). Gould et. al notes several design friction choices on Twitter’s platform to warn users against potential misinformation in some posts, as well as asking users to confirm they have read an article themselves before re-tweeting (Gould et al., 2021).
Friction can also be a tool towards meaningful reflection and intention. It is common for functionally opposing design approaches to remove key aspects of their function in order to highlight their values. This includes works in critical design, counter-functional design, slow design, reflective design, and lucid design  (Pierce and Paulos, 2014; Sengers et al., 2005; Gaver and Martin, 2000). Physical friction can also be designed to induce user reflection. For example, “Crank That" is a physical hand-crank intervention to increase friction and awareness of passive consumption in Twitter content feeds (Song et al., 2021). Users are required to continuously turn a hand crank in order to power a screen to use social media, allowing users to “question how they value various interactions on social media by transforming their usage into an embodied one in which they have to physically work for their experience” (Song et al., 2021).
@@
An effective proactive mechanism should prevent or deter unwanted behavior. Currently, only a few messaging platforms allow users to prevent screenshots altogether. [^2] https://www.safedigitalintimacy.org/state-of-the-industry StopNCII, a hashing-based system developed by multiple stakeholders including Meta, allows users to upload photos to a server before these photos are ever posted by someone else, and then block subsequent attempts by someone else to upload these photos. [^3] https://stopncii.org/ However, even then there are many drawbacks. Transformed images may evade the hash algorithm. When these images continue to be reposted, platforms may overly rely on StopNCII and invest less effort in removals. And of course, users must trust the third party–in this case, Meta. Our work aims to make the burden more equitable, placing “the work” on a potential perpetrator rather than the potential victim.
@@ 3. Formative Study
We conducted a formative interview study to develop design requirements for a technological intervention. The formative study aimed to augment prior qualitative and survey research on NCIM (Citron and Franks, 2014; Citron, 2018, 2022; Franks, 2014; Geeng et al., 2020; Powell et al., 2019; McGlynn et al., 2021; Freed et al., 2017; Qin et al., 2024). The first portion of the interview focused on participants’ existing screenshot experiences and practices. The second portion asked participants for their thoughts about using a screenshot deterrence system. In our conception of Hands-Off, we believed that a gestural interface–one that prevented viewers from using the screenshot features by occupying their hands–could be an effective deterrent. We used the formative study as an opportunity to test this idea. Thus, in the second part of the interview, we had participants explore how they would use a screenshot deterrence feature. We described the design of a gesture-based approach and solicited reflection and feedback feedback. Our university’s IRB exempted the study.
@@ 3.1. Participants
Eight participants were recruited from two undergrad courses at a large university, consisting of 4 women and 4 men. We specifically recruited participants who were able and willing to share their own or other peoples’ experiences about screenshots. We wanted to capture not only NCIM-specific experiences but also wider patterns of unauthorized screenshot sharing so that Hands-Off may be a system used for messaging broadly. We gave participants the option of either 1:1 solo interviews or paired interviews in which the participant brought a friend, who often acted as a confidante, to encourage more conversation. Each interview session took between 45 to 60 minutes. Each participant was compensated \$15 for a pair interview or \$20 for a solo interview. In the paired interview, both participants were paid and both generated information used for the study.
@@ 3.2.2. Theme 2: Perceived security of gesture-based screenshot prevention
We presented participants with a possible gesture-based approach, using hand gestures to both indicate preferences about and deter unwanted screenshots. Participants expressed a strong preference for gesture-based security features within mobile messaging applications. Their feedback shed light on the underlying motivations and practical benefits. Gesture-based security provides a tangible barrier, making it significantly harder for the recipient to share content.
@@
Participants noted that gesture-based deterrence could reduce the potential awkwardness associated with directly requesting confidentiality through written or spoken messages. This suggests that a gestural approach may enhance privacy while streamlining communication.
@@ 3.2.3. Theme 3: Gesture should depend on context
All participants believed that there should be a variety of gestures to choose from to account for various usage contexts. Specifically, context depends on three axes: the type of content sent, the relationship with the recipient, and the recipient’s location of viewing. Table 1 maps the connection between the axis and its design considerations.
@@
Table 1. Design considerations from formative study. The type of gesture matters for of content sent, the relationship between sender and recipient, and the location for viewing the content.
@@ 4.1.1. Threat model
Figure 3. Threat model of possible methods of screen duplication after content is sent. Duplicated content may be posted online repeatedly, continuing this cycle. Hands-Off contributes to deterrence against screenshotting content. Also listed is prior work and where it fits in this threat model: preventing photography((Bo et al., 2014; Zhu et al., 2017)), preventing online access((Ilia et al., 2015b; Liu et al., 2022; Hassan et al., 2017; Vishwamitra et al., 2022)), and preventing duplication((Zhang et al., 2015; Pias et al., 2022)).
The threat model for Hands-Off is unwanted photo duplication in online sharing (Bedi et al., 2013) (see Fig. 3). The system should deter general unauthorized duplication, which would take media out of its original context. We borrow from prior visual privacy systems and “assume that the adversaries are the users, the online services and third parties (e.g., external observers)” (Olteanu et al., 2018). The paradigm begins with a sender who sends a piece of original content to a recipient. From there, the recipient may download the file, screen record, screenshot, capture it using a second device themselves, or utilize an accomplice to capture on a different device.
Hands-Off goals include preventing screenshots, and deterring second device capture by the same individual. Some portions of the threat model are accounted for in prior work. In-app downloads and screen recording may be disabled by the system developer, and second device capture may be deterred with a system such as Kaleido that uses lighting patterns to disrupt outside cameras (Zhang et al., 2015). Screenshots on the device may also be prevented within the operating system of mobile devices, but this is rarely implemented in popular mobile messaging applications. As we have observed, there is no approach that can absolutely guarantee that an image not be captured and reshared. However, it is worth considering which approaches raise the cost of different types of “attacks”.
@@ 4.1.2. Design goals
Hands-Off functions by occupying both hands to prevent individuals from engaging the screenshot features on the device. As the hands are occupied, this may further deter second device capture. [^4] Clearly, this is just a deterrent. A particularly motivated attacker could position a second device, record a video, clip frames, and recover a fairly good representation. However, our analysis of themes in the formative study reveals additional specific design constraints:
@@
In addition to the explicit guidelines above, we have opted to focus on abstract gestures with relatively little cultural significance. Some gestures are culturally dependent in meaning (e.g., palms touching in a prayer gesture) whereas others have existing iconic or metaphoric roles (e.g., hands making a heart symbol). By relying on more abstract gestures, we hope that their meaning may be more flexible. This allows some degree of communicable understanding through gestures, while still allowing flexibility for how users choose to interpret them. That said, depending on context, gestures with existing meaning may be desirable for certain applications. These are possible with Hands-Off, but are left to future work.
@@ 4.2. Design space
Hands-Off gestures are shaped with two key factors in mind: preventing misuse and fitting within societal norms. Like a literary symbol, social gestures are imbued with meaning. They communicate through physicalization of language or ceremony (Kang and Tversky, 2016; Schiffrin, 1974). A wave goodbye between friends, a handshake after a negotiation, or a raised hand in a court of law are all means of communicating between parties but also to observers. They have semantic meaning (Xu et al., 2009), which may be learned, innate (Iverson and Goldin-Meadow, 1997), or inferred (e.g., hands with palms outward to represent a ‘wall’) (Müller et al., 2013). Even a gesture made in private can embody a sentiment or belief or serve a regulatory function (Rodríguez and Palacios, 2007). Using a gesture, even an abstract one, can physicalize an affirmation, whether observed by someone else or not: “I understand that the image was shared with me and is not intended for sharing with others.” The friction imposed by the physical nature of the gesture further acts to reinforce the meaning and, thus, a broader value system of consent, autonomy, and agency. It is not a button to be clicked away.
As a form of body language, the choice of gestures in interaction must take into account likely social interpretation. This is key as Hands-Off exists within two overlapping, but distinct, design spaces: bodily friction and social friction. Bodily friction is the invisible but physical restraint. The constrained space is the gap between the user when performing the necessary gesture, and the physical location of the two phone buttons that need to be pressed simultaneously to screenshot. In Hands-Off, the gap is maintained by the system’s detection of a hand gesture and the distance between reaching over and pressing the two buttons. Along the social dimension, someone using their phone in a public space will consider their social appearance in doing so. Prior work has shown that people generally want to avoid making social blunders in the presence of others, and subtle movements and movements that look and feel like everyday actions are preferred over uncommon movements (Rico and Brewster, 2010).
@@ 4.3. Use cases
Silly moment just for us. Sam is having a good time and wants to share a photo of him out drinking with his long-distance best friend Ray. However, Sam doesn’t want Ray to send this photo to anyone else. He chooses Hands-Off and selects the binoculars option to match the silliness of the photo with the gesture, and also to imply Ray should not reshare this photo. When Ray receives the message, he is prompted to perform the hand gesture and then sees the photo Sam has shared.
@@ 4.4. Implementation details
We built Hands-Off using SwiftUI for iOS devices. We chose the iPhone for its consistent button placement and hardware design. We rely on pose detection using a custom CoreML hand pose detection model trained with a custom corpus we built. We defined a multi-class neural-network classification model with 20 convolution layers for interlace, wave, frame, binoculars, and a background class for when none of the above gestures are detected. Training data consisted of 120 to 200 photos of hands performing each gesture per class. We collected the vast majority of the data ourselves by taking photos of the hands of participants. We account for high and low lighting, light and dark skin tones, and different photography angles. The remaining photos are copyright-free images of hands from the internet. The model was trained to recognize the hands through the positions of the fingertips and the locations of the joints. See Table 2 for model performance. Facial data was not used to train the model.
@@ 4.5. Hands-Off limitations
4.5.1. System limitations.
One limitation of using the CoreML model is it consistently recognized interlaced hands as one hand instead of two, which impacted precision for interlace (see Table 2). For each other hand gesture, Hands-Off required two hands detected as well as gesture detected to display the photo. However, we had to silently change this two-hands criterion for the interlace option. This limitation tells us the CoreML vision model works well for detected separated hands or poses, but may be limited for gestures when the hands stack or otherwise overlay each other.
We note several limitations in Hands-Off. First, Hands-Off is based on assumptions and modes of interactions specific to the iPhone. Screenshot deterrence is built on top of one static assumption of the double button press on the left and right sides of the phone body. The methods of screenshots can vary in other designs of mobile phones. As future work progresses, limitations may also arise in securing hand gesture-based consent in AR, VR, and XR spaces due to the possibility of certain gestures already being used as input by the specific system. Second, this study tackles just one aspect of many potential privacy concerns in the threat model. Screen recording can still be used to capture the screen if that feature is not disabled in the application. Future research in gesture-based deterrence should consider allowing users to create their own custom hand gestures. This personalization would add a uniquely intimate layer to the interaction, similar to how secret handshakes signify closeness between two people or some in-group status.
4.5.2. Accessibility limitations.
Currently, Hands-Off is limited in accommodating all abilities (e.g., people with mobility impairments and those with blindness or low vision). Feminism should acknowledge intersectionality, and that includes building for various abilities. Expanding the range of gesture-based methods to accommodate users with different abilities, needs, and various device models is crucial. For individuals with limited hand dexterity or those who may not be able to perform certain gestures, providing customizable gestures or alternative input methods, such as voice commands or facial expressions, could enhance usability. For users with visual impairments, integrating haptic feedback or audio cues to guide and confirm interactions can make the system more accessible. Compatibility with screen readers and providing clear, descriptive labels and instructions for all gestures and functions would also be important. Accessibility testing with diverse user groups may be a continuous part of the development process to identify and address any barriers that may arise. By prioritizing inclusive design, Hands-Off can better serve a broader range of users and embody principles of intersectional feminism.